OpenStudy (anonymous):

The slope of the secant line to the function f(x)=sqrt(x) between x=1 and x=16

myininaya (myininaya):

\[\frac{f(16)-f(1)}{16-1}=\frac{\sqrt{16}-\sqrt{1}}{16-1}=\frac{4-1}{15}=\frac{3}{15}\] this can be simplified a little more